OpenStudy (anonymous):

for A and B this X^3-AX^2+X-B divide on X^2+X-6 i need to find A and B..

OpenStudy (anonymous):

This is a bit challenging! If X^3-AX^2+X-B is going to divide X^2+X-6, then X^3-AX^2+X-B = (X^2+X-6)(X + ?) for some ? that we need to figure out.

OpenStudy (anonymous):

frustrating question (:

OpenStudy (anonymous):

the answers is -8 and 42

OpenStudy (anonymous):

I messed up on one thing. Now it's fixed Multiplying everything out and simplifying, you get X^3-AX^2+X-B = x^3 + (1 + ?) x^2 + (? - 6)x - 6? And from the symmetry here, we know ? - 6 = 1 and thus ? = 7

OpenStudy (anonymous):

-A = 1 + ? implies A = -8 B = 6? implies B = 42

OpenStudy (anonymous):

Normally I walk through things interactively, but this one seemed so tough, I thought I'd just put the whole thing out here. Any questions? There may be other approaches, but that's the best way I can think of. Perhaps you could do it a long division way or something.

OpenStudy (anonymous):

umm.. no questions. thanks man!

OpenStudy (anonymous):

hehe...sure, no problem.